1. What is the magnitude of the electric field E at a field point 2.0 m from a point charge q=4.0 nC?
1
Expert's answer
2021-05-31T17:08:48-0400

"E=\frac{kq}{r^2}"

Where: E is electric field

r is distance=2M

q is charge=4nC

"k = constant = 9\u00d710^{9}"

"E= \frac{9×10^9×4×10^{-9}}{2^2}=9N/C"

E=9N/C